# Mingyu Kim1∗, Pronoy Sarker1, Seungmo Kim1, Daniel J. Stilwell2, and Jorge Jimenez3

### v2

@@ I Introduction
In marine and subsurface settings, sensing quality varies over space and time due to turbidity, sediment burial, and biofouling [1, 4, 5, 6]. In aerial applications, atmospheric and illumination factors, such as cloud cover, haze, strong sunlight, rain, and smoke, likewise modulate detection performance across space and time [7, 8, 9].
@@
We model target arrivals with a Cox process so that spatial, temporal, and environmental uncertainty are represented by a random intensity. In particular, we adopt a log–Gaussian Cox process (LGCP) where logarithm of its intensity function is modelled as Gaussian process [17, 18, 19, 20]. For data-driven studies with AIS vessel traffic, we estimate the intensity using integrated nested Laplace approximation (INLA) [21, 22]. To account for false alarms, we introduce an availability function that couples detection with false-alarm behavior and filtering and quantifies the loss of effectiveness from resource drain. This function incorporates the dynamics of the external factors. We then ask when filtering helps rather than hurts and provide a verifiable condition for that regime. Furthermore, with these stochastic target and sensing models, we search for the near-optimal sensor locations by applying greedy selection, which is computationally efficient. For performance guarantee analysis, we derive a lower bound on void probability with a coverage condition that guarantees strictly better than the classical lower bound 11/e1-1/e of the optimality. Finally, we establish finite-sample robustness to explore practical sensor placement considering false alarms and filtering under a dynamical environment. When detection probabilities are learned from limited data, concentration results and robust optimization ideas [23, 24, 25] ensure that placement quality degrades in a controlled way, and results from adaptive and noisy submodularity [26, 27, 28] justify stable near-optimal performance.

### v3

@@ I Introduction
In marine and subsurface settings, sensing quality varies over space and time due to turbidity, sediment burial, and biofouling [1, 4, 5, 6]. Additional environmental effects, such as sediment properties and seafloor interaction, further influence sensing and detection, as shown in recent seabed-embedment studies [7]. In aerial applications, atmospheric and illumination factors, such as cloud cover, haze, strong sunlight, rain, and smoke, likewise modulate detection performance across space and time [8, 9, 10].
@@
Prior research provides several relevant foundations. Earlier work established Poisson-based spatial sensing models and optimal placement principles for static environments [18, 3]. Follow-on studies examined stochastic target trajectories and barrier-coverage behavior [19], improved approximation of spatial sensing performance in challenging seabed acoustic environments [20], and explored data-driven characterizations of maritime uncertainty and vessel behavior [21, 22]. Additional recent work investigated Poisson-based outlier detection in seabed sensing networks [23]. These results collectively highlight the importance of modeling spatial, temporal, and environmental variability when designing real-world sensor networks.
@@
We model target arrivals with a Cox process so that spatial, temporal, and environmental uncertainty are represented by a random intensity. In particular, we adopt a log–Gaussian Cox process (LGCP) where the logarithm of its intensity function is modelled as a Gaussian process [24, 25, 26, 27]. For data-driven studies with AIS vessel traffic, we estimate the intensity using integrated nested Laplace approximation (INLA) [28, 29].
To account for false alarms, we introduce an availability function that couples detection with false-alarm behavior and filtering and quantifies the loss of effectiveness from resource drain. This function incorporates the dynamics of the external factors. We then ask when filtering helps rather than hurts and provide a verifiable condition for that regime.
Furthermore, with these stochastic target and sensing models, we search for the near-optimal sensor locations by applying greedy selection, which is computationally efficient. For performance-guarantee analysis, we derive a lower bound on void probability with a coverage condition that guarantees strictly better than the classical lower bound 11/e1-1/e of the optimality. Finally, we establish finite-sample robustness to explore practical sensor placement considering false alarms and filtering under a dynamical environment. When detection probabilities are learned from limited data, concentration results and robust optimization ideas [30, 31, 32] ensure that placement quality degrades in a controlled way, and results from adaptive and noisy submodularity [33, 34, 35] justify stable near-optimal performance.